During 2015, a company sells 220 units of inventory. The company has the following inventory purchase transactions for 2018 (6 points)
  

Date

Transaction

Number of Units

Unit Cost

Total Cost

Jan. 1

Beginning inventory

115

$60

$6900

Sep. 8

Purchase

210

62

13020

Nov 18

Purchase

85

75

6375

410

$26,295

Calculate cost of goods sold and ending inventory for 2015

Specific ID

Sold all of beginning inventory, 25 from Nov 18 purchase, and the rest out of Sep 8 purchase.

FIFO

LIFO

AVERAGE COST

BONUS QUESTION:  At 1/1/2019 the allowance had a debit balance of 5600. Jackson Inc. wrote off $23,000 of customer accounts on Jun5, 2019. Customers repaid Johnson $18,500 of the accounts that were written off on August 20, 2019. At 12/31/2019 AR was 250,000 and estimated amount uncollectible was 15%. Jackson uses the allowance method for write offs. (3 points total)

Given the above answer the following questions

  1. What was the balance in the allowance account before adjustment( debit or credit)
  2. What is bad debt expense at 12/31/2019?
  3. What is the ending balance in the allowance account at 12/31/2019
Answer #1
Specific ID:
Cost of units sold will be as follows:
Units Cost Total cost
From beginning inventory 115 60 6900
From Nov 18. purchase 25 75 1875
From Sep 8. purchase 80 62 4960
(220-115-25)
Total 13735
Cost of goods sold=$ 13735
Cost of ending inventory:
Units Cost Total cost
From beginning inventory 0 60 0
(115-115)
From Nov 18. purchase 60 75 4500
(85-25)
From Sep 8. purchase 130 62 8060
(210-80)
Total 190 12560
Cost of ending inventory=$ 12560
FIFO:
In this method goods purchased first are sold first
Cost of units sold will be as follows:
Units Cost Total cost
From beginning inventory 115 60 6900
From Sep 8. purchase 105 62 6510
(220-115)
Total 13410
Cost of goods sold=$ 13410
Cost of ending inventory:
Units Cost Total cost
From beginning inventory 0 60 0
(115-115)
From Sep 8. purchase 105 62 6510
(210-105)
From Nov 18. purchase 85 75 6375
Total 190 12885
Cost of ending inventory=$ 12885
LIFO:
In this method goods purchased last are sold first
Cost of units sold will be as follows:
Units Cost Total cost
From Nov 18. purchase 85 75 6375
From Sep 8. purchase 135 62 8370
(220-85)
Total 220 14745
Cost of goods sold=$ 14745
Cost of ending inventory:
Units Cost Total cost
From beginning inventory 115 60 6900
From Sep 8. purchase 75 62 4650
(210-135)
From Nov 18. purchase 0 75 0
(85-85)
Total 190 11550
Cost of ending inventory=$ 11550
Average cost:
Cost of goods sold=Units sold*Average cost per unit
Average cost per unit=Total cost of goods available for sale/Total units available for sale=26295/410=$64.13 per unit
Cost of goods sold=220*64.13=$ 14109
Cost of ending inventory=Units in ending inventory*Average cost per unit
Units in ending inventory=Total units available for sale-Units sold=410-220=190
Cost of ending inventory=190*64.13=$ 12185
Bonus question:
Balance in the allowance account before adjustment:
$
Beginning balance (Debit) -5600
Add: Accounts written off now reversed 18500
12900
Less:Accounts written off 23000
Ending balance (Debit) -10100
Required balance in allowance account=Accounts receivable*15%=250000*15%=$ 37500 (Credit)
Balance in the allowance account before adjustment=10100 (Debit)
Bad debt expense=37500+10100=$ 47600
Ending balance in allowance account=$ 37500 (Credit)